1. Part A.Acetaminophen is a weak organic acid with a pKa of 9.5. What percentage of...

1.

Part A.Acetaminophen is a weak organic acid with a pKa of 9.5. What percentage of a given dose will be in the lipid soluble form at an intestinal pH of 7.5%?

Part B.Psueduoephedrine is a weak organic base with a pKa of 9.8. What will be the ratio of non-lipid soluble drug to lipid soluble drug at a stomach pH of 2.8? As the dose of psudoephedrine which you have just take moves from your stomach with a pH of 2.8 to your jejunum with a pH of 7.8, will the absorption of the drug into your blood stream increase or decrease? Do the math to prove it.
Answer #1

1.

Part A

In the case of weak acids and weak bases the ability to move from aqueous to lipid or vice versa varies with the pH of the medium.

• Henderson-Hasselbalch equation (can be use to predict the effect of pH change on ABSORPTION):

weak Acid: pH = pKa + log [A-]/[HA]

Non ionized form [HA] of a drug is lipid soluble .

Acetaminophen is a weak organic acid with a pKa of 9.5. What percentage of a given dose will be in the lipid soluble form at an intestinal pH of 7.5 :

we have,

pH = pKa + log [A-]/[HA]  

7.5 = 9.5 + log [A-]/[HA]      ===> log [A-]/[HA] = -2

[A-]/[HA] = 0.01

[A-] = [HA] * 0.01

thus, lipid soluble form :   1/1.01 *100 = 99 %

Part B.

Psueduoephedrine is a weak organic base with a pKa of 9.8. What will be the ratio of non-lipid soluble drug to lipid soluble drug at a stomach pH of 2.8?

Non ionized form [A] of a drug is lipid soluble, and ionized form protonated form will be non-lipid soluble[HA] .

we have,

pH = pKa + log [A-]/[HA]  

2.8 = 9.8 + log [A ]/[HA]      ===> log [A]/[HA] = -7

[A]/[HA] = 10-7

thus , ratio of non-lipid soluble drug to lipid soluble drug at a stomach pH of 2.8 ; [HA]/[A] = 107

As the dose of psudoephedrine which you have just take moves from your stomach with a pH of 2.8 to your jejunum with a pH of 7.8 .

here,

pH = pKa + log [A-]/[HA]  

7.8 = 9.8 + log [A ]/[HA]      ===> log [A]/[HA] = -2

[A]/[HA] =0.01

[A] = 1%

So, it can be stated that absorption of the drug into your blood stream increase